Debbie's Cookies has a return on assets of 8.1 percent and a cost of equity of 12.5 percent. What is the pretax cost of debt if the debt–equity ratio is .87? Ignore taxes.
Answer #1

Solution:-

Weight of equity = 1 / ( 1 + 0.87) = 0.5348

Weight of debt = 1 - 0.5348 = 0.4652

0.081 = 0.5348 * 0.125 + 0.4652 * pre tax cost of debt

0.081 = 0.06685 + 0.4652*pre tax cost of debt

0.01415 = 0.4624*pre tax cost of debt

pre tax cost of debt = 0.03060 or 3.06%